OpenStudy (anonymous):

Factor completely. 125−27y^3

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@Compassionate would the answer be (5-3y)(9y^2+15y+25)??

OpenStudy (compassionate):

You should recognize this as "Difference of cubes" given in most textbooks as: a^3 - b^3 = (a - b)(a^2 + ab + b^2) : In our problem you can see 125 = 5^3 and 27 = 3^3, so we have: : (5 - 3y)(25 + (5*3y) + 9y^2 which is: (5 - 3y)(25 + 15y + 9y^2)

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

wait...compassionateeee....(3wz^3 - 2s) (9w^2z^6 + 6wz^3s + 4s^2) for this equation..6wz^3s what does this mean?? 6wz^3s^3???

OpenStudy (anonymous):

@Compassionate